# Signatures of cubic gravity in the strong regime

### v1

@@ I Introduction
On an arbitrary four-dimensional spacetime, in addition to $\mathcal{P}$, there exist two other cubic densities, $\mathcal{C}$ and $\mathcal{C}^{\prime}$ (see Eq. $(2.16)$ in Ref. [3] for the full expression), which preserve the properties mentioned above and provide a rich phenomenology. For instance, in a cosmological context where spacetime is described by the Friedmann-Lemaître-Robertson-Walker metric, the (unique) combination that maintains these properties is $\mathcal{P}-8\mathcal{C}$. This theory is called Cosmological Einsteinian cubic gravity [4] and has been applied to describe the dynamics of inflation and late-time cosmological epochs [5, 6, 7, 8].111 In Ref. [2], the quartic extension of Einsteinian cubic gravity is presented, while Ref. [9] introduces and studies the quartic extension of Cosmological Einsteinian cubic gravity.
@@
For a static and spherically symmetric background in vacuum, the densities $\mathcal{C}$ and $\mathcal{C}^{\prime}$ contribute only trivially to the field equations, leaving $\mathcal{P}$ as the unique non-trivial combination. In this context, cubic gravity admits the existence of black hole solutions where the temporal and radial metric components are inversely related [10, 11], as in the Schwarzschild solution. However, the background geometry typically deviates from the Schwarzschild solutions due to the inclusion of the cubic term, resulting in a rich additional phenomenology (e.g. Refs. [12, 13, 14, 15, 16]). As this work focuses solely on spherically symmetric vacuum solutions, the densities $\mathcal{C}$ and $\mathcal{C}^{\prime}$ are not considered.
@@
Like other higher-order gravitational theories, Einsteinian cubic gravity introduces new degrees of freedom compared to general relativity. In order to prevent their propagation and ensure that the theory retains the same spectrum as general relativity, these extra modes are considered infinitely heavy (by construction). However, this model exhibits unstable modes when perturbations over a static and spherically symmetric solution are considered [17, 18]. Similar pathologies arise in the cosmological version, with the presence of at least one ghost mode and a short-time-scale tachyonic instability in the regime of small anisotropies (see Ref. [19] for details). In this sense, cubic gravity cannot be considered a complete theory but rather a perturbative correction (or effective low-energy theory) to general relativity. In other words, this model describes gravitational dynamics at low energies up to a certain energy scale, beyond which it exhibits Ostrogradski instabilities. As discussed in Refs. [17, 18, 20], for static and spherically symmetric spacetimes, the theory remains free of pathologies only within a specific regime where the Einsteinian cubic gravity is treated as an effective theory, $|\lambda\mathcal{P}/M_{\textrm{Pl}}^{3}|\ll|R|$. Here, $R$ denotes the Ricci scalar, $M_{\textrm{Pl}}$ the reduced Planck mass, and $\lambda$ the coupling constant.222 The fact that our domain satisfies $|\lambda\mathcal{P}/M_{\textrm{Pl}}^{3}|\ll|R|$ implies that all non-perturbative effects arising from the cubic term, as well as quantum effects, are excluded. When these effects come into play, the breakdown of the effective theory approach leads to the previously mentioned problems [17, 18].
@@
The paper is organized as follows: In Section II, we first introduce the field equations that govern the dynamics of a static, spherically symmetric spacetime in Einsteinian cubic gravity. Then, we provide and discuss analytical and approximate solutions to this system. A family of maximally symmetric solutions is presented in Subsection II.1, while the approximated iteratively weak coupling, asymptotic, and near-horizon solutions are presented and discussed in Subsections II.2II.3, and II.4, respectively. These solutions extend those previously considered in Refs. [10, 11, 13, 30], motivated by similar mathematical considerations. In particular, we extend the analysis to both positive and negative coupling constant and to non-asymptotically flat spacetimes expressing all approximate solutions in terms of a single integration constant. To conclude this section, we discuss the horizon properties in Subsection II.5 and the validity regimes of the iterative solutions in Subsection II.6. Possible signatures of the cubic terms, arising from their modification of light deflection, are addressed in Section III. In Subsection III.1, the main ideas behind the methodology for calculating the angular difference are examined, while the results are presented in Subsection III.2. In Subsection III.3, we discuss the modifications that appear in the effective potential for a solution with the mass of SgrA and their implications for the photon sphere position.
Finally, we conclude in Sec. IV with an overview of our results. Some complementary material is presented in the appendixes.
@@ II Static, spherically symmetric solutions
where $M_{\textrm{Pl}}:=1/\sqrt{8\pi G}$ is the reduced Planck mass, $R$ represents the Ricci scalar, and $\Lambda$ is a constant with units of inverse energy squared, potentially related to the cosmological constant, often referred to as the bare cosmological constant. (The observed value of the cosmological constant in the theory at hand is obtained as a combination of $\Lambda$ and the contributions from the cubic terms. For a detailed discussion on the cosmological constant and its various contributions, see Ref. [31].) The parameter $\lambda$ is a coupling constant with units of energy and characterizes the energy scale at which the contributions of the cubic terms
@@
The vacuum field equations (in their covariant form) are obtained by varying the action (1) with respect to the metric tensor $g^{\mu\nu}$,
@@
Here, $G_{\mu\nu}$ is the Einstein tensor and $P_{\mu\nu}$, provided in Appendix A.1, represents the modification introduced by the cubic terms Eq. (II) to the vacuum covariant equations of general relativity.
For a static, spherically symmetric solution with a single metric function $f(r)$
@@ II.1 Maximally symmetric solutions
We now show the existence of maximally symmetric solutions in the theory described by action (1). A maximally symmetric solution describes a spacetime that is homogeneous and isotropic. Such a spacetime satisfies the conditions that: $R$ is a constant, $R_{\mu\nu}=g_{\mu\nu}R/4$, and $R_{\mu\nu\gamma\rho}=R(g_{\mu\gamma}g_{\nu\rho}-g_{\nu\gamma}g_{\mu\rho})$.
@@
$0\displaystyle 2C_{0}-\frac{16r^{3}}{M_{\textrm{Pl}}^{3}}\left(\Lambda_{\text{Eff}}^{3}\lambda+\frac{M_{\textrm{Pl}}^{5}}{16}\Lambda_{\text{Eff}}-\frac{M_{\textrm{Pl}}^{5}\Lambda}{48}\right)=$ 0 . (9)
Taking the integration constant $C_{0}=0$, Eq. (9) reduces to a cubic algebraic equation for $\Lambda_{\text{Eff}}$ that admits real roots. Therefore, we obtain that a maximally symmetric solution is present in cubic gravity. Analogous to the de Sitter solution, the roots of $\Lambda_{\text{Eff}}$ act as an effective cosmological constant, which is given in terms of the bare cosmological constant $\Lambda$ and the coupling parameter $\lambda$. This modulation of the cosmological constant is known in the literature as self-tuning. Notice that the positive (negative) $\Lambda_{\text{Eff}}$ roots behave similarly to the de Sitter (anti-de Sitter) solution. A discussion of the real roots of Eq. (9) is presented in Appendix A.2.
@@ II.2 Weak coupling solution
From Eq. (II), in the limit where the cubic gravity contribution is non-dominant ($\lambda\ll 1$, weak coupling), it is natural to consider the metric function Ansatz:
@@
where $c_{1}$ and $c_{2}$ are constants with units of inverse energy and inverse energy squared, respectively. However, after evaluating this Ansatz in Eq. (II), the resulting equation becomes:
@@
and one can verify that vacuum solutions describing a black hole in empty spacetime with the structure in Eq. (10) are not possible in this gravity model. Therefore, it becomes necessary to consider an approximate black hole solution with a more general structure for the metric function:
@@
Inserting the Ansatz (12) into Eq. (II) and identifying the different orders in $\epsilon$, one arrives at a set of $m$ differential equations that can be solved iteratively for the functions $h_{n}$.333 A similar analysis can be carried out using the equation of motion (II). We refer the reader to Eq. $2.15$ in Ref. [11] for the results. In this case, a set of integration constants, $C_{m}$, appears when solving the set of differential equations iteratively. These integration constants will be related to the constant $C_{0}$ through physical properties of the object, such as its horizon radius. For the first three functions we have:
@@ II.3 Asymptotic solution
sign($\Lambda_{\text{Eff}}\Lambda$) -- $+$ or -- $+$ or -- $+$
Table 1: Summary of real-space solutions for the effectively constant $\Lambda_{\text{Eff}}$. The regions where the sign of the cosmological constant $\Lambda$ matches the sign of effective constant $\Lambda_{\text{Eff}}$ are free from ghost instabilities. Notice that for $\lambda\geq-M_{\textrm{Pl}}^{5}/(12\Lambda^{2})$, there exists at least one branch of solutions where this occurs.
@@
where $a_{n}$ and $b_{n+1}$ are constants coefficients. The terms proportional to $r^{n}$ include the expected de Sitter asymptotic behavior, while those proportional to $r^{-(n+1)}$ capture the asymptotic gravitational potential due to a massive source.
Evaluating Eq. (14) in Eq. (II) and considering the first ten terms (which require $m=11$ due to the differentiations involved), we find that, as expected, all coefficients $a_{n}$ are zero, with the exception of $a_{0}=1$ and $a_{2}=\Lambda_{\text{Eff}}$. The $b_{n}$ coefficients are determined iteratively, leading to the asymptotic solution:
@@
The ellipsis indicates the remaining terms, shown in the Eq. A.4 in Appendix A.4 and the effective cosmological constant $\Lambda_{\text{Eff}}$ is determined from the real root of the algebraic equation:
@@
$\triangle>0$: For $\lambda>0$ or $\lambda<-\frac{M_{\textrm{Pl}}^{5}}{12\Lambda^{2}}$ we have that $\triangle>0$ and Eq. (16) has one real root given by:
@@
It is straightforward to prove from Eq. (17) that the sign of the real root is opposite to that of the cosmological constant $\Lambda$ when $\lambda<-\frac{M_{\textrm{Pl}}^{5}}{12\Lambda^{2}}$ and matches it when $\lambda>0$ ensuring that this branch remains free from ghosts [11].
$\triangle<0$: When $-\frac{M_{\textrm{Pl}}^{5}}{12\Lambda^{2}}<\lambda<0$, then $\triangle<0$ and all roots of Eq. (16) are real and are given by,
@@
$\triangle=0$: Finally, when $\lambda=-\frac{M_{\textrm{Pl}}^{5}}{12\Lambda^{2}}$, then $\triangle=0$, all roots are real, and two of them are identical:
$\Lambda_{\text{Eff}}=\frac{3\Lambda}{2}\;\text{(doubled root)},\quad\Lambda_{\text{Eff}}=-3\Lambda.$ (19)
@@ II.4 Near Horizon solution
In this section, we proceed to identify the power series solution of Eq. (II) near a horizon. Considering the Ansatz (4) for the metric, the black hole horizon is defined as a surface of radius $r=r_{h}$ where $f(r_{h})=0$. Furthermore, we assume $f^{\prime}(r_{h})\geq 0$ at the (non-cosmological) horizon. This is justified since we are interested in the exterior black hole horizon, where the metric function $f(r)$ is expected to change sign from positive (outside the horizon) to negative (inside the horizon). Notice, however, that we are allowing for $f(r)$ to have a minimum at $r_{h}$, since this is also compatible with $f(r_{h})=0$ and $f(r>r_{h})>0$. According to these constraints, we consider the Ansatz:
@@
Considering the solution branch that yields a finite value for $f_{\text{nhs}}^{\prime}(r_{h})$ in the limit $\lambda=0$, one can use the first-order results to obtain an equation for the horizon radius in terms of the constants $C_{0},\lambda$, and $\Lambda$,
@@ II.5 Horizons properties
Figure 1: Discrete solutions space of Eq. (24): The main plot displays a set of solutions $(r_{h},f^{\prime}(r_{h}))$ for a range of $\lambda$ values, indicated by a color gradient. The black line denotes the solutions obtained in II.4, while the red markers are some illustrative numerical solutions. In particular, the red triangle marker indicates the solution for $\lambda=0.05$, reported in Fig. 2. The secondary (right) plot illustrates the existence of a cosmological horizon across the entire solution space. A detailed discussion of this figure is provided in the main text. The results are expressed in the units defined in (27).
@@
Focusing on the exterior (non-cosmological) black hole horizon, we observe that at $r_{h}$ the coefficients of $f^{\prime\prime}$ in Eq. (II) vanish, resulting in a cubic equation for the first derivative at the horizon:
@@
Now, when only $\Lambda=0$, Eq. (24) reduces to a quadratic equation for $r_{h}$, whose solutions can be written (after some manipulations) as:
@@
From Eq. (II.5), one can infer that there exist solutions whose only horizon is equal to the Schwarzschild radius ($r_{s}={2C_{0}}/{M_{\textrm{Pl}}^{2}}$), independently of the values of $\lambda$, these solutions have $f^{\prime}(r_{h})=0$. Notice that the same conclusion can be obtained from Eq. (24) with $f^{\prime}(r_{h})=0$. The fact that Eq. (II) is well-behaved when $f\to 0$ and $f^{\prime}\to 0$ simultaneously suggest that the metric function is smooth at the horizon. In a recent work, Ref. [35], the author presents a novel class of solutions featuring naked singularities with these characteristics. In particular, our results explain why, for all values of $\lambda$, the position of the critical horizon remains equal to the Schwarzschild radius. Continuing with the analysis for $\Lambda=0$ and $f^{\prime}(r_{h})>0$, let us discuss the particularities of the cases $\lambda<0$ and $\lambda>0$:
$\lambda<0$: The expression inside the brackets in Eq. (II.5) is positive for the branch with the upper signs and negative for the other branch. (Notice that the square root is greater than one.) It can be proven that there exists only one real positive solution $r_{h}$ which corresponds to the upper signs in Eq. (II.5). [^4] The lower signs always lead to $r_{h}<0$, this can be shown by taking Eq. (24) with $\Lambda=0$ and rewriting it as a quadratic for $r_{h}$ where the coefficient of the quadratic term is positive and the coefficients of the linear and zero-order terms are negative. Thus, from the general quadratic formula, it is easy to verify that the solution with the lower sign is always negative. In this case, the right-hand side of Eq. (II.5) is less than one, indicating a horizon radius greater than the Schwarzschild radius, i.e. $r_{h}(\lambda<0;\Lambda=0)>r_{h}(\lambda=0;\Lambda=0)$. This result is consistent with the findings of Ref. [10].
$\lambda>0$: In this case, both roots $r_{h}$ of Eq. (II.5) are positive in the region determined by:
@@
In this region, the expression inside the brackets in Eq. (II.5) is always positive, which implies that the right-hand side is greater than one. This indicates that both solutions for $r_{h}$ are smaller than the Schwarzschild radius, i.e. $r_{h}(\lambda>0;\Lambda=0)<r_{h}(\lambda=0;\Lambda=0)$, for both the upper and lower signs in Eq. (II.5).

### v2

@@ I Introduction
On an arbitrary four-dimensional spacetime, in addition to $\mathcal{P}$, there exist two other cubic densities, $\mathcal{C}$ and $\mathcal{C}^{\prime}$ (see Eq. (2.16)(2.16)( 2.16 ) in Ref. [3] for the full expression), which preserve the properties mentioned above and provide a rich phenomenology. For instance, in a cosmological context where spacetime is described by the Friedmann-Lemaître-Robertson-Walker metric, the (unique) combination that maintains these properties is $\mathcal{P}-8\mathcal{C}$. This theory is called Cosmological Einsteinian cubic gravity [4] and has been applied to describe the dynamics of inflation and late-time cosmological epochs [5, 6, 7, 8, 9].111 In Ref. [2], the quartic extension of Einsteinian cubic gravity is presented, while Ref. [10] introduces and studies the quartic extension of Cosmological Einsteinian cubic gravity.
@@
For a static and spherically symmetric background in vacuum, the densities $\mathcal{C}$ and $\mathcal{C}^{\prime}$ contribute only trivially to the field equations, leaving $\mathcal{P}$ as the unique non-trivial combination. In this context, cubic gravity admits the existence of black hole solutions where the temporal and radial metric components are inversely related [11, 12], as in the Schwarzschild solution. However, the background geometry typically deviates from the Schwarzschild solution due to the inclusion of the cubic term, resulting in a rich additional phenomenology (e.g. Refs. [13, 14, 15, 16, 17]). As this work focuses solely on spherically symmetric vacuum solutions, the densities $\mathcal{C}$ and $\mathcal{C}^{\prime}$ are not considered.
@@
Like other higher-order gravitational theories, Einsteinian cubic gravity introduces new degrees of freedom compared to general relativity. In order to prevent their propagation and ensure that the theory retains the same spectrum as general relativity, these extra modes are considered infinitely heavy (by construction). However, this model exhibits unstable modes when perturbations over a static and spherically symmetric solution are considered [18, 19]. Similar pathologies arise in the cosmological version, with the presence of at least one ghost mode and a short-time-scale tachyonic instability in the regime of small anisotropies (see Ref. [20] for details). In this sense, cubic gravity cannot be considered a complete theory but rather a perturbative correction (or effective low-energy theory) to general relativity. In other words, this model describes gravitational dynamics at low energies up to a certain energy scale, beyond which it exhibits Ostrogradsky instabilities. As discussed in Refs. [18, 19, 21], for static and spherically symmetric spacetimes, the theory remains free of pathologies only within a specific regime where the Einsteinian cubic gravity is treated as an effective theory, $|\lambda\mathcal{P}/M_{\textrm{Pl}}^{3}|\ll|R|$. Here, $R$ denotes the Ricci scalar, $M_{\textrm{Pl}}$ the reduced Planck mass, and $\lambda$ the coupling constant.222 The fact that our domain satisfies $|\lambda\mathcal{P}/M_{\textrm{Pl}}^{3}|\ll|R|$ implies that all non-perturbative effects arising from the cubic term, as well as quantum effects, are excluded. When these effects come into play, the breakdown of the effective theory approach leads to the previously mentioned problems [18, 19].
@@
The paper is organized as follows: In Section II, we first introduce the field equations that govern the dynamics of a static, spherically symmetric spacetime in Einsteinian cubic gravity. Then, we provide and discuss analytical and approximate solutions to this system. A family of maximally symmetric solutions is presented in Subsection II.1, while the approximated iterative weak coupling, asymptotic, and near-horizon solutions are presented and discussed in Subsections II.2II.3, and II.4, respectively. These solutions extend those previously considered in Refs. [11, 12, 14, 31], motivated by similar mathematical considerations. In particular, we extend the analysis to both positive and negative coupling constant and to non-asymptotically flat spacetimes expressing all approximate solutions in terms of a single integration constant. To conclude this section, we discuss the horizon properties in Subsection II.5 and the validity regimes of the iterative solutions in Subsection II.6. Possible signatures of the cubic terms, arising from their modification of light deflection, are addressed in Section III. In Subsection III.1, the main ideas behind the methodology for calculating the angular difference are examined, while the results are presented in Subsection III.2. In Subsection III.3, we discuss the modifications that appear in the effective potential for a solution with the mass of SgrA and their implications for the photon sphere position.
Finally, we conclude in Section IV with an overview of our results. Some complementary material is presented in the appendixes.
@@ II Static, spherically symmetric solutions
where $M_{\textrm{Pl}}:=1/\sqrt{8\pi G}$ is the reduced Planck mass, $R$ represents the Ricci scalar, and $\Lambda$ is a constant with units of energy squared, potentially related to the cosmological constant, often referred to as the bare cosmological constant. (The observed value of the cosmological constant in the theory at hand is obtained as a combination of $\Lambda$ and the contributions from the cubic terms. For a detailed discussion on the cosmological constant and its various contributions, see Ref. [32].) The parameter $\lambda$ is a coupling constant with units of energy and characterizes the energy scale at which the contributions of the cubic terms
@@
The vacuum field equations (in their covariant form) are obtained by varying the action (1) with respect to the metric tensor $g^{\mu\nu}$, as follows:
@@
Here, $G_{\mu\nu}$ is the Einstein tensor and $P_{\mu\nu}$, provided in Appendix A.1, represents the modification introduced by the cubic terms in Eq. (II) to the vacuum covariant equations of general relativity.
For a static, spherically symmetric solution with a single metric function $f(r)$,
@@ II.1 Maximally symmetric solutions
We now show the existence of maximally symmetric solutions in the theory described by action (1). A maximally symmetric solution describes a spacetime that is homogeneous and isotropic. Such a spacetime satisfies the conditions that $R$ is a constant, $R_{\mu\nu}=g_{\mu\nu}R/4$, and $R_{\mu\nu\gamma\rho}=R(g_{\mu\gamma}g_{\nu\rho}-g_{\nu\gamma}g_{\mu\rho})$.
@@
$2C_{0}-\frac{16r^{3}}{M_{\textrm{Pl}}^{3}}\left(\Lambda_{\text{Eff}}^{3}\lambda+\frac{M_{\textrm{Pl}}^{5}}{16}\Lambda_{\text{Eff}}-\frac{M_{\textrm{Pl}}^{5}\Lambda}{48}\right)=0$ (9)
Equation (9) can be solved for all $r$ only if $C_{0}=0$, consistent with the fact that $C_{0}$ represents the ADM mass of the solution, which is not present in the Ansatz (8). In this case, Eq. (9) reduces to a cubic algebraic equation for $\Lambda_{\text{Eff}}$ that admits real roots. Therefore, we obtain that a maximally symmetric solution is present in cubic gravity. Analogous to the de Sitter solution, the roots of $\Lambda_{\text{Eff}}$ act as an effective cosmological constant, which is given in terms of the bare cosmological constant $\Lambda$ and the coupling parameter $\lambda$. This modulation of the cosmological constant is known in the literature as self-tuning. Notice that the positive (negative) $\Lambda_{\text{Eff}}$ roots behave similarly to the de Sitter (anti-de Sitter) solution. A discussion of the real roots of Eq. (9) is presented in Appendix A.2.
@@ II.2 Weak coupling solution
From Eq. (II), in the limit where the cubic gravity contribution is non-dominant ($\lambda\ll 1$, weak coupling), it is natural to consider the metric function Ansatz
@@
where $c_{1}$ and $c_{2}$ are constants with units of inverse energy and inverse energy squared, respectively. However, after evaluating this Ansatz in Eq. (II), the resulting equation becomes
@@
and one can verify that vacuum solutions describing a black hole in empty spacetime with the structure in Eq. (10) are not possible in this gravity model. Therefore, it becomes necessary to consider an approximate black hole solution with a more general structure for the metric function
@@
Inserting the Ansatz (12) into Eq. (II) and identifying the different orders in $\epsilon$, one arrives at a set of $m$ differential equations that can be solved iteratively for the functions $h_{n}$.333 A similar analysis can be carried out using the equation of motion (II). We refer the reader to Eq. (2.152.152.15) in Ref. [12] for the results. In this case, a set of integration constants, $C_{m}$, appears when solving the set of differential equations iteratively. These integration constants will be related to the constant $C_{0}$ through physical properties of the object, such as its horizon radius. For the first three functions, we have
@@ II.3 Asymptotic solution
[…]
Table 1: Summary of real-space solutions for the effective constant $\Lambda_{\text{Eff}}$. The regions where the sign of the cosmological constant $\Lambda$ matches the sign of effective constant $\Lambda_{\text{Eff}}$ are free from ghost instabilities. Notice that for $\lambda\geq-M_{\textrm{Pl}}^{5}/(12\Lambda^{2})$, there exists at least one branch of solutions where this occurs.
@@
where $a_{n}$ and $b_{n+1}$ are constant coefficients. The terms proportional to $r^{n}$ include the expected de Sitter asymptotic behavior, while those proportional to $r^{-(n+1)}$ capture the asymptotic gravitational potential due to a massive source.
Evaluating Eq. (14) in Eq. (II) and considering the first ten terms (which require $m=11$ due to the differentiations involved), we find that, as expected, all coefficients $a_{n}$ are zero, with the exception of $a_{0}=1$ and $a_{2}=\Lambda_{\text{Eff}}$. The $b_{n}$ coefficients are determined iteratively, leading to the asymptotic solution
@@
The ellipsis indicates the remaining terms, shown in Eq. (A.4) in Appendix A.4 and the effective cosmological constant $\Lambda_{\text{Eff}}$ is determined from the real root of the algebraic equation
@@
>0$\triangle$>0△ > 0: for $\lambda>0$ or $\lambda<-\frac{M_{\textrm{Pl}}^{5}}{12\Lambda^{2}}$ we have that >0$\triangle$>0△ > 0 and Eq. (16) has one real root given by:
@@
It is straightforward to prove from Eq. (17) that the sign of the real root is opposite to that of the cosmological constant $\Lambda$ when $\lambda<-\frac{M_{\textrm{Pl}}^{5}}{12\Lambda^{2}}$ and matches it when $\lambda>0$ ensuring that this branch remains free of ghosts [12].
<0$\triangle$<0△ < 0: when $-\frac{M_{\textrm{Pl}}^{5}}{12\Lambda^{2}}<\lambda<0$, then <0$\triangle$<0△ < 0 and all roots of Eq. (16) are real and are given by,
@@
=0$\triangle$=0△ = 0: finally, when $\lambda=-\frac{M_{\textrm{Pl}}^{5}}{12\Lambda^{2}}$, then =0$\triangle$=0△ = 0, all roots are real, and two of them are identical:
$\Lambda_{\text{Eff}}=\frac{3\Lambda}{2}\;\text{(double root)},\quad\Lambda_{\text{Eff}}=-3\Lambda.$ (19)
@@ II.4 Near Horizon solution
In this section, we proceed to identify the power series solution of Eq. (II) near a horizon. Considering the Ansatz (4) for the metric, the black hole horizon is defined as a surface of radius $r=r_{h}$ where $f(r_{h})=0$. Furthermore, we assume $f^{\prime}(r_{h})\geq 0$ at the (non-cosmological) horizon. This is justified, since we are interested in the exterior black hole horizon, where the metric function $f(r)$ is expected to change sign from positive (outside the horizon) to negative (inside the horizon). Notice, however, that we are allowing for $f(r)$ to have a minimum at $r_{h}$, since this is also compatible with $f(r_{h})=0$ and $f(r>r_{h})>0$. According to these constraints, we consider the Ansatz
@@
Considering the solution branch that yields a finite value for $f_{\text{nhs}}^{\prime}(r_{h})$ in the limit $\lambda=0$, one can use the first-order results to obtain an equation for the horizon radius in terms of the constants $C_{0},\lambda$, and $\Lambda$, as follows:
@@ II.5 Horizons properties
Figure 1: Discrete solution space of Eq. (24): the main plot displays a set of solutions $(r_{h},f^{\prime}(r_{h}))$ for a range of $\lambda$ values, indicated by a color gradient. The black line denotes the solutions obtained in II.4, while the red markers are some illustrative numerical solutions. In particular, the red triangle marker indicates the solution for $\lambda=0.05$, reported in Fig. 2. The secondary (right) plot illustrates the existence of a cosmological horizon across the entire solution space. A detailed discussion of this figure is provided in the main text. The results are expressed in the units defined in (27).
@@
Focusing on the exterior (non-cosmological) black hole horizon, we observe that at $r_{h}$ the coefficients of $f^{\prime\prime}$ in Eq. (II) vanish, resulting in a cubic equation for the first derivative at the horizon, as follows:
@@
Now, when only $\Lambda=0$, Eq. (24) reduces to a quadratic equation for $r_{h}$, whose solutions can be written (after some manipulations) as
@@
From Eq. (II.5), one can infer that there exist solutions whose only horizon is equal to the Schwarzschild radius ($r_{s}={2C_{0}}/{M_{\textrm{Pl}}^{2}}$), independently of the values of $\lambda$, these solutions have $f^{\prime}(r_{h})=0$. Notice that the same conclusion can be obtained from Eq. (24) with $f^{\prime}(r_{h})=0$. The fact that Eq. (II) is well behaved when $f\to 0$ and $f^{\prime}\to 0$ simultaneously suggest that the metric function is smooth at the horizon. In a recent work, Ref. [36], the author presents a novel class of solutions featuring naked singularities with these characteristics. In particular, our results explain why, for all values of $\lambda$, the position of the critical horizon remains equal to the Schwarzschild radius. Continuing with the analysis for $\Lambda=0$ and $f^{\prime}(r_{h})>0$, let us discuss the particularities of the cases $\lambda<0$ and $\lambda>0$:
$\lambda<0$: the expression inside the brackets in Eq. (II.5) is positive for the branch with the upper signs and negative for the other branch. (Notice that the square root is greater than 1.) It can be proven that there exists only one real positive solution $r_{h}$ which corresponds to the upper signs in Eq. (II.5). [^4] The lower signs always lead to $r_{h}<0$, this can be shown by taking Eq. (24) with $\Lambda=0$ and rewriting it as a quadratic for $r_{h}$ where the coefficient of the quadratic term is positive and the coefficients of the linear and zero-order terms are negative. Thus, from the general quadratic formula, it is easy to verify that the solution with the lower sign is always negative. In this case, the right-hand side of Eq. (II.5) is less than one, indicating a horizon radius greater than the Schwarzschild radius, i.e. $r_{h}(\lambda<0;\Lambda=0)>r_{h}(\lambda=0;\Lambda=0)$. This result is consistent with the findings of Ref. [11].
$\lambda>0$: in this case, both roots $r_{h}$ of Eq. (II.5) are positive in the region determined by
@@
In this region, the expression inside the brackets in Eq. (II.5) is always positive, which implies that the right-hand side is greater than 1. This indicates that both solutions for $r_{h}$ are smaller than the Schwarzschild radius, i.e. $r_{h}(\lambda>0;\Lambda=0)<r_{h}(\lambda=0;\Lambda=0)$, for both the upper and lower signs in Eq. (II.5).